OpenStudy (anonymous):

What equation would be perpendicular to y=x+2

OpenStudy (anonymous):

do you know what coefficient m should like in perpendicular lines

OpenStudy (anonymous):

m2=-1/m1

OpenStudy (anonymous):

since in original line the slope is 1, then the line that is perpendicular will have slope -1.

OpenStudy (anonymous):

@Bassfishingmaster

OpenStudy (anonymous):



OpenStudy (anonymous):

these lines are perpendicular

OpenStudy (anonymous):

As to your question, the lines that are perpedicular to y=x+2 will be y=-x, y=-x+7, y=-x-76 and other.